O yes QUESTION 2 A wave has period of 20 s. The frequency is Hz. QUESTION 3 A wave has frequency of 200 Hz. The period is s QUESTION 4 A rope wave travels at 20 m /s with a frequency of 4 Hz. The wavelength ism. Click Save and Submit to save and submit. Click Save All Answers to save all answers.
/shsu blackboard.com/webapps/assessment/take/aunch.jsp?course assessment id- 220022 18course id- 118791.18 v Question Completion Status: QUESTIONS A rope wave has a wavelength of 2 m and a frequency of 3 Hz. The velocity of the wave ism/s. QUESTION 6 Are rope waves transverse or longitudinal? transverse O longitudinal QUESTION 7 Are water waves, as when dropping a stone in a pond, transverse or longitudinal? O transverse O longitudinal QUESTION 8 Click Save and Submit to save and submit.Click Save All Answers to save all answers
d.com/webapps/assessment/take/launchjsp? course assessment id 220022 18course id 118791 1&content.id 3535273 18step ¥ Question Completion Status: QUESTION 8 Are sound waves transverse or longitudinal? O transverse O longitudinal QUESTION9 The intensity of a 60 W light bulb 2 m away isW m2 QUESTION 10 The intensity of a 60 W light bulb 6 m away isWm2 QUESTION 11 What fraction of the energy emitted from the sun is absorbed by the Earth? Treat the sun like a light bulb 93 million miles away. Als assume the Earth only absorbs the energy associated with an area of approximately 50 million square miles. (This is roughly the are disc with the same radius as the Earth.) Click Save and Submit to save and submit. Cliek Save All Ansvers to save all answers
hsu blackboard.com/webapps/assessment/take/launchjspcourse assessment id- 220022 18course id- 118791 1contentd 3535273.18step-nul Question Completion Status: QUESTION 12 By what factor would the intensity of a wave have to increase to add 3 dB? QUESTION 13 By what factor would the intensity of a wave have to increase to add 10 dB? QUESTION 14 By what factor would the intensity of a wave have to increase to add 100 dB QUESTION 15 Medical devices exist that can combine waves at the locations of kidney stones to destroy them. Are these devices using constructive or destructive interference? O constructive O destructive Click Save and Submit to save and submit. Click Save All Ansuvers to save all answers
course.assessment.ld 220022 1&course.id 118791 1content.id3535273 1step- ¥ Question Completion Status: Medical devices exist that can combine waves at the locations of kidney stones to destroy them. Are these devices using constructiv destructive interference? O constructive O destructive QUESTION 16 Anti-noise headphones can generate sound waves that cancel out unwanted background noise. Are these devices using constructive destructive interference? O constructive O destructive QUESTION 17 is measured while the ambulance is stationary, moving away from the receiver, and, moving towards An ambulance siren frequency stationary receiver. The three frequencies measured are 896 Hz, 960 H z and 1034 Hz. Which frequency was measured while the ambulance was moving towards the receiver? O 896 Hz O 960 Hz O 1034 Hz anawers Click Save and Submit to save and submit. Click Save All Anscers to save all
Answer #1

2.The frequency f=1/T=1/20=0.05 Hz.

3.The period=1/f=1/200=0.005 s.

4.The speed of the wave=v=flambda . v=20 m/s and f=4 Hz. Therefore lambda =5 m.

5.The velocity of the wave=v=flambda =3Hz*2m=6 m/s.

6.Rope waves are transverse.

7.Water waves are also transverse but are two dimensional unlike rope waves.

8 Sound waves are longitudinal.

9.The intensity of the bulb 2 m away=1.2VVn.

10.Intensity=I=I=I_{0}/4pi r^{2}=60 W/144pi =0.13 W/m^{2}.

11.The fraction of energy absorbed=(50 million square miles)/(93 million miles)2=6*10-9 .

12.We have the relation = (10dB)log1 / 10 . Now D+ 3 = (10dB)logT / 10 .

Subtracting the 2 equations we obtain logi, / I = 0.3 i.e I'=2I.

13.In the same way as in the previous case I = 101.

14 Here 1, 10101

15. The devices are using constructive interference,

16.The devices are using destructive interference.

17.The frequency measured is greatest when the source is moving towards the receiver. Therefore the correct answer is the the third one i.e 1034 Hz.